A box with a square base and no top is to be made from a square piece of cardboard by cutting 3 in. squares from each corner and folding up the sides. The box is to hold 2352 in^3 (inches cubed). How big a piece of cardboard is needed?
I think I understand but I can't get the right answer. Please help :)

Answer by stanbon(75887)

A box with a square base and no top is to be made from a square piece of cardboard by cutting 3 in. squares from each corner and folding up the sides. The box is to hold 2352 in^3 (inches cubed). How big a piece of cardboard is needed?
--------------
Volume = 3(x-6)^2 = 2352
(x-6)^2 = 786
x -6 = 28.04
x = 34.04
Ans: 34.04^2 = 1158.72 sq in
--------------
Cheers,
Stan H.
==============